OpenStudy (anonymous):

can someone thoroughly explain this please? limit as x approaches 81 f(x)=(sqrtx)-9/x-81

OpenStudy (anonymous):

wait...\[{\lim_{x \rightarrow 81}}\frac{ \sqrt{x}-9 }{ x-81 }\] or \[{\lim_{x \rightarrow 81}}\frac{ \sqrt{x-9} }{ x-81 }\]?

OpenStudy (anonymous):

nvm you mean the first one right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Since substituting 81 in for x will result in an undefined result ...0/0

OpenStudy (anonymous):

So you have to use another method...Have you learned l'Hopital's Theorem yet?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

hmmmm....

OpenStudy (anonymous):

I got this wrong on a quiz so is there another way to do it?

OpenStudy (anonymous):

Well what did you put on your quiz?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

And the correct answer was 1/18, right? Or don't you know?

OpenStudy (anonymous):

Yes, that answer was correct. But, I don't know how to get that answer.

OpenStudy (anonymous):

Ok ... Wow well. Without using l'Hopital's I am at a loss. What course are you in?

OpenStudy (anonymous):

AP Calc. is I'Hopital"s simple? Maybe my teacher assumed that everyone knew it already

OpenStudy (anonymous):

OK then you can handle it and that might be the case. Here it is in a nutshell.

OpenStudy (anonymous):

If a limit results in 0/0 or infinity/infinity or -infinity/-infinity you can use l'Hoptial's

OpenStudy (anonymous):

It says:\[\lim_{x \rightarrow c}\frac{ f(x) }{ g(x)}=\lim_{x \rightarrow c}\frac{ f'(x) }{ g'(x) }\]

OpenStudy (anonymous):

So this means that if you can't get the answer by subbing in c for x into the expression b/c it results in 0/0, etc. then take the derivative of the numerator and of the denominator and try again.

OpenStudy (anonymous):

So for this problem...\[\lim_{x \rightarrow 81}\frac{ \sqrt{x} -9}{ x-81 }=\frac{ \sqrt{81} -9}{ 81-81 }=\frac{ 0 }{ 0 }\] you can use l'Hopital's

OpenStudy (anonymous):

That means...

OpenStudy (anonymous):

\[\lim_{x \rightarrow 81}\frac{ \sqrt{x}-9 }{ x-81 }=\lim_{x \rightarrow 81}\frac{ \frac{ 1 }{ 2\sqrt{x} } }{ 1 }=\lim_{x \rightarrow 81}\frac{ 1 }{ 2\sqrt{x} } =\frac{ 1 }{ 2\sqrt{81} }=\frac{ 1 }{18 }\]

OpenStudy (anonymous):

How is that @Mynameee ?

OpenStudy (anonymous):

Where did the 1/2 come from?

OpenStudy (anonymous):

The derivative of \[\sqrt{x}\] is

OpenStudy (anonymous):

if you havent covered lhopitals rule you can solve it without it by multiply like this : \[\frac{ \sqrt{x}-9 }{ x-81 }*\frac{ \sqrt{x}+9 }{ \sqrt{x}+9 }\] then see what you get , some things will cancel out :)

OpenStudy (anonymous):

\[\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

Oh! That looks familiar @litchlani

OpenStudy (anonymous):

Good call @litchlani but l'Hopital is so much fun to say!

OpenStudy (anonymous):

i know lhopital makes things always so easy but if they havent covered it their prof. probably wanted them to do that

OpenStudy (anonymous):

Totally agree ... I am just talking about how much fun it is to say l'Hopital.